OpenStudy (anonymous):

Find P(A) P(A and B) = 1/12, P(B given A) = 2/3

OpenStudy (anonymous):

P(A and B) = P(A) * P(B given A) 1/12 = P(A) * 2/3 P(A) = 1/8